A 20 kW motor lifts a load at a speed of 0.4 m/s. What is the maximum load it can lift at this speed
1
Expert's answer
2018-08-23T11:19:08-0400
The formula for the power is
P=Fv.
The maximum load it can lift at this speed is
F=P/v=20000/0.4=50 kN.
Answer: 50 kN.